OpenStudy (anonymous):

1. In the parallelogram above, if the measure of W is 74°, what is the measure of X? 74° 69° 106° 286°

OpenStudy (aravindg):

where is the figure?

OpenStudy (anonymous):

https://www11.studyisland.com/userfiles/tx6par.gif

OpenStudy (anonymous):

there it is

OpenStudy (aravindg):

well do you know the properties of angles of a parallelogram?

OpenStudy (anonymous):

not rlly..u kno wat it is

OpenStudy (anonymous):

i mean the answer

OpenStudy (aravindg):

yep X+W=Y+Z=180 degree

OpenStudy (aravindg):

now can you get it ?

OpenStudy (anonymous):

no 180 isnt on the answer list

OpenStudy (aravindg):

X+W=Y+Z=180

OpenStudy (aravindg):

you know W find X

OpenStudy (anonymous):

ohh 106

OpenStudy (aravindg):

;)